OpenStudy (anonymous):

Simplify the given expression: 4/3-2i

OpenStudy (campbell_st):

multiply by the conjugate pair... 3 + 2i and remember i^2 = -1 the process will give a fractional with a rational denominator \[\frac{4}{3 - 2i} \times \frac{3 + 2i}{3 + 2i}\] hope it helps